OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. (2 points) 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (anonymous):

@jim_thompson5910 @seiga @Mertsj @blarghhonk8

OpenStudy (anonymous):

P1: 2=1(1+1) Pk: 2 + 4 + 6 + . . . + 2k = k(k+1) Pk+1: 2 + 4 + 6 + . . . + 2(k+1) = (k+1)[(k+1)+1)]

OpenStudy (anonymous):

since thy are asking for the statements that means i can leave it like this? i don't have to prove it? @myko

OpenStudy (anonymous):

you can leave it like this. Pn is for any n: for example n=1, or n=k, or n=k+1

OpenStudy (anonymous):

ok. THANK YOU SOOOO MUCH!!!! @myko

OpenStudy (anonymous):

yw